# Learning Content-enhanced Mask Transformer for Domain Generalized Urban-Scene Segmentation

### 09 Dec 2023 v4

@@ Abstract
Domain-generalized urban-scene semantic segmentation (USSS) aims to learn generalized semantic predictions across diverse urban-scene styles. Unlike domain gap challenges, USSS is unique in that the semantic categories are often similar in different urban scenes, while the styles can vary significantly due to changes in urban landscapes, weather conditions, lighting, and other factors. Existing approaches typically rely on convolutional neural networks (CNNs) to learn the content of urban scenes.
@@
Extensive experiments conducted on various domain-generalized urban-scene segmentation datasets demonstrate that the proposed CMFormer significantly outperforms existing CNN-based methods for domain-generalized semantic segmentation, achieving improvements of up to 14.00% in terms of mIoU (mean intersection over union). The source code is publicly available at https://github.com/BiQiWHU/CMFormer.
@@ Introduction
Urban-scene semantic segmentation (USSS) is a challenging problem because of the large scene variations due to changing landscape, weather, and lighting conditions (Sakaridis, Dai, and Van Gool 2021; Mirza et al. 2022; Diaz-Ruiz et al. 2022; Chen et al. 2022). Unreliable USSS can pose a significant risk to road users. Nevertheless, a segmentation model trained on a specific dataset cannot encompass all urban scenes across the globe. As a result, the segmentation model is prone to encountering unfamiliar urban scenes during the inference stage. Hence, domain generalization is essential for robust USSS (Pan et al. 2018; Huang et al. 2019a; Choi et al. 2021), where a segmentation model can effectively extrapolate its performance to urban scenes that it hasn’t encountered before (Fig. 1). In contrast to common domain generalization, domain generalized USSS requires special attention because the domain gap is mainly caused by large style variations whereas changes in semantics largely remain consistent (e.g. as shown in the example in Fig. 2).
@@
The proposed Content-enhanced mask attention (CSMA) mechanism can be integrated into existing mask-level segmentation Transformer in a learnable fashion. It consists of three key steps, namely, exploiting high-resolution properties, exploiting low-resolution properties, and content-enhanced fusion. Besides, it can also be seamlessly adapted to multi-resolution features. A novel Content-enhanced Mask TransFormer (CMFormer) is proposed for domain-generalized USSS.
@@
  • Extensive experiments show a large performance improvement over existing SOTA by upto 14.00% mIoU, and HGFormer by upto 1.7% mIoU.
@@
Domain Generalization
has been studied on no task-specific scenarios in the field of both machine learning and computer vision. Hu et al. (Hu and Lee 2022) proposed a framework for image retrieval in an unsupervised setting. Zhou et al. (Zhou et al. 2020) proposed a framework to generalize to new homogeneous domains. Qiao et al. (Qiao, Zhao, and Peng 2020) and Peng et al. (Peng, Qiao, and Zhao 2022) proposed to learn domain generalization from a single source domain. Many other methods have also been proposed (Zhao et al. 2020; Mahajan, Tople, and Sharma 2021; Wang et al. 2020; Chattopadhyay, Balaji, and Hoffman 2020; Segu, Tonioni, and Tombari 2023).
Domain Generalized Semantic Segmentation
can be regarded as the exploration of unsupervised domain adaption segmentation (Peng et al. 2022, 2021; Yue et al. 2019a), but requires a larger generalization capability of a model on a variety of target domains. Existing methods focus on the generalization of in-the-wild (Piva, de Geus, and Dubbelman 2023), scribble (Tjio et al. 2022) and multi-source images (Kim et al. 2022; Lambert et al. 2020), where substantial alterations can occur in both the content and style.
Domain Generalized USSS
focuses on the generalization of driving-scenes (Cordts et al. 2016; Yu et al. 2018; Neuhold et al. 2017; Ros et al. 2016; Richter et al. 2016). These methods use either a normalization transformation (e.g., IBN (Pan et al. 2018), IN (Huang et al. 2019a), SAN (Peng et al. 2022)) or whitening transformation (e.g., IW (Pan et al. 2019), ISW (Choi et al. 2021), DIRL (Xu et al. 2022), SAW (Peng et al. 2022)) on the training domain, to enable the model to generalize better on the target domains. Other advanced methods for domain generalization in segmentation typically rely on external images to incorporate more diverse styles (Lee et al. 2022; Zhao et al. 2022; Zhong et al. 2022; Li et al. 2023), and leverage content consistency across multi-scale features (Yue et al. 2019b). To the best of our knowledge, all of these methods are based on CNN.
Mask Transformer for Semantic Segmentation uses the queries in the Transformer decoder to learn the masks, e.g., Segmenter (Strudel et al. 2021), MaskFormer (Cheng, Schwing, and Kirillov 2021). More recently, Mask2Former (Cheng et al. 2022) further simplified the pipeline of MaskFormer and achieves better performance.
@@
Similar content. In Fig. 2a, samples from BDD100K and from GTA5 have same scene categories, namely, tree, cars, road and building. Without loss of generality, samples from all the five USSS datasets contain the same 19 scene-object categories. These scene-object categories all commonly exist in urban-scenes, either seen or unseen during training.
Varied styles. In Fig. 2a, we present an example sourced from BDD100K, showcasing a rainy, daytime scene from a European street, among other factors. In contrast, the GTA5 sample embodies a clear, nighttime scene from an American street, along with other attributes. Without compromising the breadth of representation, samples across all five USSS datasets exhibit an even broader spectrum of stylistic variations. These variations span weather conditions such as clear, rainy, and snowy scenarios, while also encompassing different times of day—CityScapes predominantly features daytime scenes, whereas GTA5 portrays scenes from both day and night. Regrettably, despite the pronounced diversity in style, this extensive variation may still not encompass the entirety of potential conditions in urban environments.
@@
Figure 3: In the domain-generalized USSS setting, within the content-style space, samples from various domains tend to cluster closely along the content dimension while displaying dispersion along the style dimension. (b) An optimal generalized semantic segmentation scenario would involve uniform distribution of styles while maintaining content stability (as indicated by the brown bounding box).
@@
Recent work has shown that mask-level segmentation Transformer (e.g., Mask2Former) (Ding et al. 2023) is a scalable learner for domain generalized semantic segmentation. Unfortunately, we empirically observed that, a mask-level representation is better at representing content, but more sensitive to style variations (similar to Fig. 3a); its low-resolution counterpart, on the contrary, is less capable to represent content, but more robust to the style variations (similar to the style dimensions in Fig. 3b).
Overall, the mask representation and its down-sampled counterpart shows complementary properties when handling samples from different domains. Thus, it is natural to jointly leverage both mask representation and its down-sampled counterparts, so as to at the same time stabilizing the content and be insensitive to the style variation.
@@
In contrast, the proposed method intends to leverage the content representation ability of mask-level features (content dimension of Fig. 3a) and the style handling ability of its down-sampled counterpart (style dimension of Fig. 3b), so as to realize the same learning objective.
@@
It learns the query features as the segmentation masks by introducing a mask attention matrix based on the self-attention mechanism. Let $\mathbf{F}_{l}$, $\mathbf{X}_{l}$ denote the image features from the image decoder, and the features of the $l^{th}$ layer in a Transformer decoder, respectively. When $l=0$, $\mathbf{X}_{0}$ refers to the input query features of the Transformer decoder.
The key $\mathbf{K}_{l}$ and value $\mathbf{V}_{l}$ on $\mathbf{F}_{l-1}$ are computed by linear transformations $f_{K}$ and $f_{V}$, respectively. And the query $\mathbf{Q}_{l}$ on $\mathbf{X}_{l-1}$ is computed by linear transformation $f_{Q}$. Then, the query feature $\mathbf{X}_{l}$ is computed by
@@
Highlighted within the green block in Figure 4, our empirical observations reveal that the high-resolution mask representation exhibits the following characteristics: 1) greater proficiency in content representation, and 2) reduced susceptibility to domain variation. However, achieving uniform mixing of samples from four domains presents a challenge.
To leverage the properties from high-resolution mask representations, we use the self-attention mechanism to exploit the amplified content representation from $\mathbf{X}_{l}$. Let $\mathbf{Q}_{\mathbf{X}_{l}}$, $\mathbf{V}_{\mathbf{X}_{l}}$ and $\mathbf{K}_{\mathbf{X}_{l}}$ denote its query, value and key. Then, the self-attention is computed as
@@
To this end, we fuse both representations $\tilde{\mathbf{X}}_{l}$ and $\tilde{\mathbf{X}}_{l}^{d}$ in a simple and straight-forward way. The fused feature $\mathbf{X}_{l}^{final}$ serves as the final output of the $l^{th}$ Transformer decoder, and it is computed as follows:
@@
Existing domain generalized USSS methods are included for comparison, namely, IBN (Pan et al. 2018), IW (Pan et al. 2019), Iternorm (Huang et al. 2019b), DRPC (Yue et al. 2019b), ISW (Choi et al. 2021), GTR (Peng et al. 2021), DIRL (Xu et al. 2022), SHADE (Zhao et al. 2022), SAW (Peng et al. 2022), WildNet (Lee et al. 2022), AdvStyle (Zhong et al. 2022), SPC (Huang et al. 2023), HGFormer (Ding et al. 2023).
@@
Method Proc. & Year Trained on GTA5 (G)
@@
IBN ECCV2018 33.85 32.30 37.75 27.90
IW CVPR2019 29.91 27.48 29.71 27.61
Iternorm CVPR2019 31.81 32.70 33.88 27.07
DRPC ICCV2019 37.42 32.14 34.12 28.06
ISW CVPR2021 36.58 35.20 40.33 28.30
GTR TIP2021 37.53 33.75 34.52 28.17
DIRL AAAI2022 41.04 39.15 41.60 -
SHADE ECCV2022 44.65 39.28 43.34 -
SAW CVPR2022 39.75 37.34 41.86 30.79
WildNet CVPR2022 44.62 38.42 46.09 31.34
AdvStyle NeurIPS2022 39.62 35.54 37.00 -
SPC CVPR2023 44.10 40.46 45.51 -
HGFormer CVPR2023 - - - -
CMFormer (Ours) 2023 55.31 49.91 60.09 43.80
@@
SYNTHIA Source Domain Table 2 reports the performance. The proposed CMFormer shows a 5.67%, 8.73% and 11.49% mIoU performance gain against the best CNN based methods, respectively. However, on the BBD-100K (B) dataset, the semantic-aware whitening (SAW) method (Peng et al. 2022) outperforms the proposed CMFormer by 1.80% mIoU. Nevertheless, the proposed CMFormer still outperforms the rest state-of-the-art methods. The performance gain of the proposed CMFormer when trained on SYNTHIA dataset is not as significant as it is trained on CityScapes or GTA5 dataset. The explanation may be that the SYNTHIA dataset has much fewer samples than GTA5 dataset, i.e., 9400 v.s. 24966, and a transformer may be under-trained.
Method Proc. & Year Trained on SYNTHIA (S)
@@
IBN ECCV2018 32.04 30.57 32.16 26.90
IW CVPR2019 28.16 27.12 26.31 26.51
Iternorm CVPR2019 - - - -
DRPC ICCV2019 35.65 31.53 32.74 28.75
ISW CVPR2021 35.83 31.62 30.84 27.68
GTR TIP2021 36.84 32.02 32.89 28.02
DIRL AAAI2022 - - - -
SHADE ECCV2022 - - - -
SAW CVPR2022 38.92 35.24 34.52 29.16
WildNet CVPR2022 - - - -
AdvStyle NeurIPS2022 37.59 27.45 31.76 -
SPC CVPR2023 - - - -
HGFormer CVPR2023 - - - -
CMFormer (Ours) 2023 44.59 33.44 43.25 40.65
@@
CityScapes Source Domain Table 3 reports the performance. As HGFormer only reports one decimal results (Ding et al. 2023), we also report one decimal results when compared with them. The proposed CMFormer (with Swin-Base backbone) shows a performance gain of 6.32%, 10.43%, 9.50% and 12.11% mIoU on the B, M, G and S dataset against the state-of-the-art CNN based method. As BDD100K dataset contains many nigh-time urban-street images, it is particularly challenging for existing domain generalized USSS methods. Still, a performance gain of 6.32% is observed by the proposed CMFormer.
@@
From Clear to Adverse Conditions The results and discussions are provided in the supplementary material.
[…]
Figure 6: Unseen domain segmentation prediction of existing CNN based domain generalized semantic segmentation methods (IBN (Pan et al. 2018), IW (Pan et al. 2019), ISW (Choi et al. 2021), SAW (Peng et al. 2022)) and the proposed CMFormer under the C $\rightarrow$ adverse domain setting. More prediction is in the supplementary material.
Ablation Studies
@@
Some generalized segmentation results on the C $\rightarrow$ B, M, G, S setting (in Fig. 5) and on the C $\rightarrow$ adverse domain setting (in Fig. 6). Compared with the CNN based methods, the proposed CMFormer shows a better segmentation prediction, especially in terms of the completeness of objects.

### 17 Dec 2023 v5

@@ Abstract
Domain-generalized urban-scene semantic segmentation (USSS) aims to learn generalized semantic predictions across diverse urban-scene styles. Unlike generic domain gap challenges, USSS is unique in that the semantic categories are often similar in different urban scenes, while the styles can vary significantly due to changes in urban landscapes, weather conditions, lighting, and other factors. Existing approaches typically rely on convolutional neural networks (CNNs) to learn the content of urban scenes.
@@
Extensive experiments conducted on various domain-generalized urban-scene segmentation datasets demonstrate that the proposed CMFormer significantly outperforms existing CNN-based methods by up to 14.0% mIoU and the contemporary HGFormer by up to 1.7% mIoU. The source code is publicly available at https://github.com/BiQiWHU/CMFormer.
@@ Introduction
Urban-scene semantic segmentation (USSS) is a challenging problem because of the large scene variations due to changing landscape, weather, and lighting conditions (Sakaridis, Dai, and Van Gool 2021; Mirza et al. 2022; Bi, You, and Gevers 2023; Chen et al. 2022). Unreliable USSS can pose a significant risk to road users. Nevertheless, a segmentation model trained on a specific dataset cannot encompass all urban scenes across the globe. As a result, the segmentation model is prone to encountering unfamiliar urban scenes during the inference stage. Hence, domain generalization is essential for robust USSS (Pan et al. 2018; Huang et al. 2019a; Choi et al. 2021), where a segmentation model can effectively extrapolate its performance to urban scenes that it hasn’t encountered before (Fig. 1). In contrast to common domain generalization, domain generalized USSS requires special attention because the domain gap is mainly caused by large style variations whereas changes in semantics largely remain consistent (example in Fig. 2).
@@
The proposed content-enhanced mask attention (CMA) mechanism can be integrated into existing mask-level segmentation Transformer in a learnable fashion. It consists of three key steps, namely, exploiting high-resolution properties, exploiting low-resolution properties, and content-enhanced fusion. Besides, it can also be seamlessly adapted to multi-resolution features. A novel Content-enhanced Mask TransFormer (CMFormer) is proposed for domain-generalized USSS.
@@
  • Extensive experiments show a large performance improvement over existing SOTA by up to 14.0% mIoU, and HGFormer by up to 1.7% mIoU.
@@
Domain Generalization has been studied on no task-specific scenarios in the field of both machine learning and computer vision. Hu et al. (Hu and Lee 2022) proposed a framework for image retrieval in an unsupervised setting. Zhou et al. (Zhou et al. 2020) proposed a framework to generalize to new homogeneous domains. Qiao et al. (Qiao, Zhao, and Peng 2020) and Peng et al. (Peng, Qiao, and Zhao 2022) proposed to learn domain generalization from a single source domain. Many other methods have also been proposed (Zhao et al. 2020; Mahajan, Tople, and Sharma 2021; Wang et al. 2020; Chattopadhyay, Balaji, and Hoffman 2020; Segu, Tonioni, and Tombari 2023).
Domain Generalized Semantic Segmentation is more practical than conventional semantic segmentation (Pan et al. 2022; Ji et al. 2021; Li et al. 2021; Ji et al. 2022; Zhou, Yi, and Bi 2021; Ye et al. 2021), which focuses on the generalization of a segmentation model on unseen target domains. Existing methods focus on the generalization of in-the-wild (Piva, de Geus, and Dubbelman 2023), scribble (Tjio et al. 2022) and multi-source images (Kim et al. 2022; Lambert et al. 2020), where substantial alterations can occur in both the content and style.
Domain Generalized USSS focuses on the generalization of driving-scenes (Cordts et al. 2016; Yu et al. 2018; Neuhold et al. 2017; Ros et al. 2016; Richter et al. 2016). These methods use either normalization transformation (e.g., IBN (Pan et al. 2018), IN (Huang et al. 2019a), SAN (Peng et al. 2022)) or whitening transformation (e.g., IW (Pan et al. 2019), ISW (Choi et al. 2021), DIRL (Xu et al. 2022), SAW (Peng et al. 2022)) on the training domain, to enable the model to generalize better on the target domains. Other advanced methods for domain generalization in segmentation typically rely on external images to incorporate more diverse styles (Lee et al. 2022; Zhao et al. 2022; Zhong et al. 2022; Li et al. 2023), and leverage content consistency across multi-scale features (Yue et al. 2019). To the best of our knowledge, all of these methods are based on CNN.
Mask Transformer for Semantic Segmentation uses the queries in the Transformer decoder to learn the masks, e.g., Segmenter (Strudel et al. 2021), MaskFormer (Cheng, Schwing, and Kirillov 2021). More recently, Mask2Former (Cheng et al. 2022) further simplifies the pipeline of MaskFormer and achieves better performance.
@@
Figure 3: (a) In the domain-generalized USSS setting, within the content-style space, samples from various domains tend to cluster closely along the content dimension while displaying dispersion along the style dimension. (b) An optimal generalized semantic segmentation scenario would involve uniform distribution of styles while maintaining content stability (as indicated by the brown bounding box).
@@
Recent work has shown that mask-level segmentation Transformer (e.g., Mask2Former) (Ding et al. 2023) is a scalable learner for domain generalized semantic segmentation. However, we empirically observe that, a mask-level representation is better at representing content, but more sensitive to style variations (similar to Fig. 3a); its low-resolution counterpart, on the contrary, is less capable to represent content, but more robust to the style variations (similar to the style dimensions in Fig. 3b).
Overall, the mask representation and its down-sampled counterpart shows complementary properties when handling samples from different domains. Thus, it is natural to jointly leverage both mask representation and its down-sampled counterparts, so as to at the same time stabilize the content and be insensitive to the style variation.
@@
In contrast, the proposed method intends to leverage the content representation ability of mask-level features and the style handling ability of its down-sampled counterpart, so as to realize the aforementioned learning objective.
@@
It learns the query features as the segmentation masks by introducing a mask attention matrix based on the self-attention mechanism. Let $\mathbf{F}_{l}$ and $\mathbf{X}_{l}$ denote the image features from the image decoder and the features of the $l^{th}$ layer in a Transformer decoder, respectively. When $l=0$, $\mathbf{X}_{0}$ refers to the input query features of the Transformer decoder.
The key $\mathbf{K}_{l}$ and value $\mathbf{V}_{l}$ on $\mathbf{F}_{l-1}$ are computed by linear transformations $f_{K}$ and $f_{V}$, respectively. Similarly, the query $\mathbf{Q}_{l}$ on $\mathbf{X}_{l-1}$ is computed by linear transformation $f_{Q}$. Then, the query feature $\mathbf{X}_{l}$ is computed by
@@
Highlighted within the green block in Figure 4, our empirical observations reveal that the high-resolution mask representation exhibits the following characteristics: 1) greater proficiency in content representation, and 2) reduced susceptibility to domain variation. Achieving uniform mixing of samples from four domains presents a challenge.
To leverage the properties from high-resolution mask representations, we use the self-attention mechanism to exploit the amplified content representation from $\mathbf{X}_{l}$. Let $\mathbf{Q}_{\mathbf{X}_{l}}$, $\mathbf{V}_{\mathbf{X}_{l}}$ and $\mathbf{K}_{\mathbf{X}_{l}}$ denote its query, value and key, and $d_{k}$ denotes their dimension. Then, the self-attention is computed as
@@
To this end, we fuse both representations $\tilde{\mathbf{X}}_{l}$ and $\tilde{\mathbf{X}}_{l}^{d}$ in a simple and straight-forward way. The fused feature $\mathbf{X}_{l}^{final}$ serves as the final output of the $l^{th}$ Transformer decoder, and it is computed as
@@
Existing domain generalized USSS methods are included for comparison, namely, IBN (Pan et al. 2018), IW (Pan et al. 2019), Iternorm (Huang et al. 2019b), DRPC (Yue et al. 2019), ISW (Choi et al. 2021), GTR (Peng et al. 2021), DIRL (Xu et al. 2022), SHADE (Zhao et al. 2022), SAW (Peng et al. 2022), WildNet (Lee et al. 2022), AdvStyle (Zhong et al. 2022), SPC (Huang et al. 2023), and HGFormer (Ding et al. 2023).
@@
Method Trained on GTA5 (G)
@@
IBN (Pan et al. 2018) 33.85 32.30 37.75 27.90
IW (Pan et al. 2019) 29.91 27.48 29.71 27.61
Iternorm (Huang et al. 2019b) 31.81 32.70 33.88 27.07
DRPC (Yue et al. 2019) 37.42 32.14 34.12 28.06
ISW (Choi et al. 2021) 36.58 35.20 40.33 28.30
GTR (Peng et al. 2021) 37.53 33.75 34.52 28.17
DIRL (Xu et al. 2022) 41.04 39.15 41.60 -
SHADE (Zhao et al. 2022) 44.65 39.28 43.34 -
SAW (Peng et al. 2022) 39.75 37.34 41.86 30.79
WildNet (Lee et al. 2022) 44.62 38.42 46.09 31.34
AdvStyle (Zhong et al. 2022) 39.62 35.54 37.00 -
SPC (Huang et al. 2023) 44.10 40.46 45.51 -
HGFormer (Ding et al. 2023) - - - -
CMFormer (Ours) 55.31 49.91 60.09 43.80
@@
SYNTHIA Source Domain Table 2 reports the performance. The proposed CMFormer shows a 5.67%, 8.73% and 11.49% mIoU performance gain against the best CNN based methods, respectively. However, on the BBD-100K (B) dataset, the semantic-aware whitening (SAW) method (Peng et al. 2022) outperforms the proposed CMFormer by 1.80% mIoU. Nevertheless, the proposed CMFormer still outperforms the rest methods. The performance gain of the proposed CMFormer when trained on SYNTHIA dataset is not as significant as it is trained on CityScapes or GTA5 dataset. The explanation may be that the SYNTHIA dataset has much fewer samples than GTA5 dataset, i.e., 9400 v.s. 24966, and a transformer may be under-trained.
Method Trained on SYNTHIA (S)
@@
IBN (Pan et al. 2018) 32.04 30.57 32.16 26.90
IW (Pan et al. 2019) 28.16 27.12 26.31 26.51
Iternorm (Huang et al. 2019b) - - - -
DRPC (Yue et al. 2019) 35.65 31.53 32.74 28.75
ISW (Choi et al. 2021) 35.83 31.62 30.84 27.68
GTR (Peng et al. 2021) 36.84 32.02 32.89 28.02
DIRL (Xu et al. 2022) - - - -
SHADE (Zhao et al. 2022) - - - -
SAW (Peng et al. 2022) 38.92 35.24 34.52 29.16
WildNet (Lee et al. 2022) - - - -
AdvStyle (Zhong et al. 2022) 37.59 27.45 31.76 -
SPC (Huang et al. 2023) - - - -
HGFormer (Ding et al. 2023) - - - -
CMFormer (Ours) 44.59 33.44 43.25 40.65
@@
CityScapes Source Domain Table 3 reports the performance. As HGFormer only reports one decimal results (Ding et al. 2023), we also report one decimal results when compared with it. The proposed CMFormer (with Swin-Base backbone) shows a performance gain of 6.32%, 10.43%, 9.50% and 12.11% mIoU on the B, M, G and S dataset against the state-of-the-art CNN based method. As BDD100K dataset contains many nigh-time urban-street images, it is particularly challenging for existing domain generalized USSS methods. Still, a performance gain of 6.32% is observed by the proposed CMFormer.
@@
From Clear to Adverse Conditions we further validate the proposed CMFormer’s performance on the adverse conditions dataset with correspondance (ACDC) (Sakaridis, Dai, and Van Gool 2021). We set the fog, night, rain and snow as four different unseen domains, and directly use the model pre-trained on CityScapes for inference. The results are shown in Table 5. It significantly outperforms existing domain generalized segmentation methods (Pan et al. 2018; Huang et al. 2019a; Pan et al. 2019; Choi et al. 2021; Li et al. 2023) by up to 10.3%, 0.5%, 11.6%, 11.1% on the fog, night, rain and snow domains, respectively.
Method Backbone Trained on Cityscapes (C)
$\rightarrow$ Fog $\rightarrow$ Night $\rightarrow$ Rain $\rightarrow$ Snow mean
IBN ResNet-50 63.8 21.2 50.4 49.6 43.7
Iternorm ResNet-50 63.3 23.8 50.1 49.9 45.3
IW ResNet-50 62.4 21.8 52.4 47.6 46.6
ISW ResNet-50 64.3 24.3 56.0 49.8 48.1
ISSA MiT-B3 67.5 33.2 55.9 53.2 52.5
CMFormer (Ours) Swin-B 77.8 33.7 67.6 64.3 60.1
@@
Figure 5: Unseen domain segmentation prediction of existing CNN based domain generalized semantic segmentation methods (IBN (Pan et al. 2018), IW (Pan et al. 2019), ISW (Choi et al. 2021), SAW (Peng et al. 2022)) and the proposed CMFormer under the C $\rightarrow$ B, M, G, S setting.
Figure 6: Unseen domain segmentation prediction of existing CNN based domain generalized semantic segmentation methods (IBN (Pan et al. 2018), IW (Pan et al. 2019), ISW (Choi et al. 2021), SAW (Peng et al. 2022)) and the proposed CMFormer under the C $\rightarrow$ adverse domain setting.
Ablation Studies
[…]
@@
Some segmentation results on the C $\rightarrow$ B, M, G, S setting and C $\rightarrow$ adverse domain setting are visualized in Fig. 5 and 6. Compared with the CNN based methods, the proposed CMFormer shows a better segmentation prediction, especially in terms of the completeness of objects.